# CriticBench: Evaluating Large Language Models as Critic

### v3

@@ Abstract
Critique ability are crucial in the scalable oversight and self-improvement of Large Language Models (LLMs). While many recent studies explore the critique ability of LLMs to judge and refine flaws in generations, how to comprehensively and reliably measure the critique abilities of LLMs is under-explored. This paper introduces CriticBench, a novel benchmark designed to comprehensively and reliably evaluate four key critique ability dimensions of LLMs: feedback, comparison, refinement and meta-feedback. CriticBench encompasses nine diverse tasks, each assessing the LLMs’ ability to critique responses at varying levels of quality granularity. Our extensive evaluations of open-source and closed-source LLMs reveal intriguing relationships between the critique ability and tasks, response qualities, and model scales. Datasets, resources and evaluation toolkit for CriticBench will be publicly released at https://github.com/open-compass/CriticBench.
@@ 1 Introduction
Critique ability, i.e. the ability to judge and refine flaws, is the key to self-improvement Yuan et al. (2024); Xu et al. (2023) of Large Language Models (LLMs). Powerful critique ability of LLMs also facilitates a more robust framework, i.e., scalable oversight Bowman et al. (2022); Saunders et al. (2022), for ensuring the AI systems growing in scale and capability remain aligned with human-desired outcomes and ethical standards.
Figure 1: Comparison between CriticBench and existing benchmarks for critique evaluation. More details about the comparison are in Appendix D.
While numerous studies Gou et al. (2024); Pan et al. (2023); Yuan et al. (2024); Xu et al. (2023) have explored the critique abilities to enhance the performance of LLMs on some specific tasks, a reliable and comprehensive framework for assessing the critique capabilities is under-explored. Most existing works Wang et al. (2023c); Li et al. (2024); Cui et al. (2023) mainly evaluate a singular aspect of critique ability, limited in the critique ability dimensions, application tasks and response qualities. Besides, they frequently utilize GPT-4 OpenAI (2023) for direct evaluation of critique quality, not to mention that GPT-4 is not entirely reliable Wang et al. (2023c); Zhang et al. (2024). These issues significantly impede the in-depth analysis and further improvement of critique capabilities.
To address the above problems, we introduce a diverse and novel benchmark, CriticBench, aiming to comprehensively and reliably measure the critique capabilities of LLMs. Specifically, to ensure the comprehensiveness of the benchmark, CriticBench evaluates the critique ability of LLMs under nine diverse and common task scenarios, including NLP tasks (summary, translation, and question answering), reasoning (math with program-of-thought Chen et al. (2023a) and chain-of-thought Wei et al. (2023)), coding (code with and without execution [^1] We hereinafter refer to “code w/. execution” as “CodeExec” and “code w/o. execution” as “CodeNE”. Liu et al. (2023a)) and alignment (general chat and harmlessness). We collect model responses of varying quality levels to be critiqued using various open-source and closed-source LLMs of different scales and capabilities.
The measurement of critique ability is formulated into four dimensions: feedback, correction, comparison, and meta-feedback (i.e., feedback of feedback), completely covering the capability requirements of the self-improvement Yuan et al. (2024) and scalable oversight Saunders et al. (2022). Since both the scalar-based and textual formats of critique are commonly used in these scenarios, CriticBench evaluate the critiques in both formats, equipped with objective Fu et al. (2023); Li et al. (2023) and subjective Li et al. (2024); Wang et al. (2023c) evaluations, respectively.
@@
To improve the reliability of subjective scoring for textual critiques, we employ high-quality human-annotated critiques as references to assist GPT-4, which has been proven effective in enhancing the quality of GPT-4 assessments Sun et al. (2024a); Kim et al. (2024). We contribute a human-in-the-loop pipeline Liu et al. (2023b) to efficiently annotate critiques, first generated by GPT-4 and then reviewed and refined by human experts.
Overall, CriticBench exhibits significant advantages over previous benchmarks on critique evaluation (Figure 1), showing great diversity in response quality granularity, critique formats, critique dimensions, and data size, allowing deeper analysis of the LLMs’ critique capabilities. Extensive evaluations of numerous LLMs reveal several intriguing phenomena as below:
  • The critique difficulty varies with tasks, where responses of mathematics and coding tasks are challenging for feedback and comparison.
@@
  • The critique difficulty correlates with the critique dimensions, where meta-feedback is more challenging than feedback.
  • The quality of critiques is inversely related to the response qualities to be critiqued, whereas high-quality responses are more difficult for LLMs to critique sufficiently.
  • The critique ability obeys the scaling law, i.e., LLMs with grander scale and general capabilities demonstrate better critique abilities.
There is also a promising trend that some open-source LLMs Bai et al. (2023); Team (2023); DeepSeek-AI (2024) are approaching state-of-the-art closed-source LLMs in critique capabilities.
@@
These phenomena promote an in-depth understanding of LLM’s critique ability, and we hope the discoveries could spur future research in this field.
##### Task Input and Response $\boldsymbol{R_{a}},\boldsymbol{R_{b}}$ $\boldsymbol{I}$: Please give me a recipe for pizza. $\boldsymbol{R_{a}}$: Here is a recipe for pizza: … $\boldsymbol{R_{b}}$: I don’t know any recipe for pizza … ##### Single-feedback $\boldsymbol{F_{s}(I,R_{b})}$ for $\boldsymbol{R_{b}}$ Generated response fails to fulfill user’s query … Following suggestions could be considered … Score for Response $\boldsymbol{R_{B}}$: 1/10 ##### Correction $\boldsymbol{CR(I,R_{b},F_{s})}$ with $F_{s}$ Here is the revision for response B: Of course! To make a pizza, the following ingredients are needed: … The steps for cooking are as follows: … ##### Comparison-based feedback $\boldsymbol{F_{c}(I,R_{a},R_{b})}$ Response A is more helpful than B … On the contrary, response B only claims it cannot solve this input … Preference Response: A ##### Meta-feedback $\boldsymbol{F_{s}(I,R_{b},F_{s})}$ Score for Feedback $\boldsymbol{F_{s}}$: 8/10 The feedback point out the errors in response B … It also provide suggestions …
Figure 2: Cases of key concepts. Red and brown represent scalar-based and textual critiques, respectively.
@@ 2 Preliminaries
We first formally define the key concepts and their corresponding notions in CriticBench. Figure 2 shows a specific case to understand these concepts.
@@
represent the user queries and generations produced by LLMs.
@@
aims to analyze and refine the generated responses. Formally, this paper studies the critique capabilities in four dimensions: (1) single-feedback or feedback $F_{s}$ involves textual analysis and a quality score, also noted as $F_{s}(I,R)$. Good feedback should not only find flaws but also provide helpful suggestions for refinement Saunders et al. (2022); (2) correction or refinement $CR$ aims to revise responses with or without feedback. Previous evaluations Li et al. (2024); Cui et al. (2023) overlook this dimension, although it is an inevitable step when letting the model improve itself Bai et al. (2022b); (3) comparison or comparison-based feedback $F_{c}$ contains a textual critique and a preference label for a pair of responses $(R_{a},R_{b})$. Such an ability is commonly used in the current RLAIF pipeline that relies on LLMs to label the preferences of responses Bai et al. (2022b); (4) Meta-feedback $F_{s}(I,R,F_{s})$, i.e., the feedback of feedback itself Saunders et al. (2022), involves a rating score reflecting the quality of $F_{s}$ and corresponding textual analysis, which is a higher-level critique dimension. Such an ability is necessary to improve the critique ability Yuan et al. (2024).
3 CriticBench Construction
Figure 3: The data construction pipeline for our proposed CriticBench. Step 1: $9$ tasks and numerous LLMs are prepared. Step 2: LLMs are employed to generate responses, which are then meticulously reviewed by human experts. Step 3: Critiques are generated by LLMs with strong critique ability, and human experts annotate them.
@@ 3 CriticBench Construction
Given the challenge of crafting scalar-based and textual critiques from scratch, we construct CriticBench using a human-in-the-loop data construction pipeline as shown in Figure 3.
3.1 Instruction Collection
@@
Instructions for $9$ distinct tasks are collected to evaluate critique capabilities comprehensively (Step 1 in Figure 3). Specifically, the benchmark includes three representative classical language tasks: summary Stiennon et al. (2020), translation Specia et al. (2020), and question-answering Mihaylov et al. (2018). Since a popular application of LLMs is to serve as a chatbot, where alignment is important to ensure the safe application of LLMs, we collect instructions from general chat scenarios Li et al. (2023) and harmlessness cases Bai et al. (2022b) to evaluate the LLMs’ critique ability for alignment. Furthermore, the reasoning and code capabilities are also fundamental for augmenting LLMs as agents Wang et al. (2023a), another important and promising application of LLMs. Thus, we also collect instructions for math reasoning with chain-of-thought and program-of-thought, and coding with and without execution results. To ensure the difficulty of CriticBench, we only collect coding and math reasoning questions that some 70B LLMs cannot correctly answer. For each task, we collect around 100 instructions from the test sets of some widely-used benchmark datasets to ensure the instruction quality and avoid data contamination. More details are in Appendix A.
3.2 Response Generation
For each collected $I$ in each task, LLMs of different scales and capabilities (e.g., Vicuna Chiang et al. (2023) and WizardLM Xu et al. (2024)) are employed to generate responses with flaws, which naturally form responses of various qualities (Step 2 (a) in Figure 3). The complete list of LLMs are in Appendix B. To identify the quality of these responses efficiently, GPT-4 is utilized to initially assign quality ratings ranging from $1$ to $7$ (Step 2 (b) in Figure 3) then let human annotators meticulously review and adjust these scores, which are used in the objective evaluation in the feedback dimension (Section 4.1). Subsequently, three responses with distinct quality differences for each $I$ are chosen based on their human-varified quality scores, including low-, medium-, and high-quality responses (noted as $R_{\rm low},R_{\rm med},R_{\rm high}$, respectively).
Moreover, golden or correct responses are collected for each task input $I$, which are proven challenging for critiques Wang et al. (2023c); Zhang et al. (2024). We use GPT-4 to generate correct responses using ground-truth rationales or codes as hints for coding and mathematical tasks. Since executions leak information about response quality, correct responses are not collected for the CodeExec task. In tasks beyond coding and mathematics, GPT-4 is prompted to refine its past generations, given its feedback during multiple turns, and the last revision is collected as golden response.
@@ 3.2 Response Generation
After sampling task inputs and their corresponding responses $R_{\rm low},R_{\rm med},R_{\rm high}$, we collect the feedback to be evaluated for the meta-feedback dimension by utilizing $4$ LLMs that are known powerful for critiques (Step 3 (d) in Figure 3): (1) GPT-4; (2) GPT-3.5-turbo; (3) Auto-J-13B Li et al. (2024); (4) UltraCM-13B Cui et al. (2023).
The average annotated scores for collected responses with varying qualities on $9$ tasks can be found in Appendix C.
3.3 Reference Critique Generation

### v4

# CriticEval: Evaluate Large Language Model as Critic

@@ Abstract
Critique ability, i.e., the capability of Large Language Models (LLMs) to identify and rectify flaws in responses, is crucial for their applications in self-improvement and scalable oversight. While numerous studies have been proposed to evaluate critique ability of LLMs, their comprehensiveness and reliability are still limited. To overcome this problem, we introduce CriticEval, a novel benchmark designed to comprehensively and reliably evaluate critique ability of LLMs. Specifically, to ensure the comprehensiveness, CriticEval evaluates critique ability from four dimensions across nine diverse task scenarios. It evaluates both scalar-valued and textual critiques, targeting responses of varying quality. To ensure the reliability, a large number of critiques are annotated to serve as references, enabling GPT-4 to evaluate textual critiques reliably. Extensive evaluations of open-source and closed-source LLMs first validate the reliability of evaluation in CriticEval. Then, experimental results demonstrate the promising potential of open-source LLMs, the effectiveness of critique datasets and several intriguing relationships between the critique ability and some critical factors, including task types, response qualities and critique dimensions. Datasets and evaluation toolkit for CriticEval will be publicly released.
@@ 1 Introduction
Critique ability is crucial for the self-improvement (yuan2024selfrewarding, ; xu2023reasons, ) of LLMs, as it enables the effective analysis and correction of flaws in responses (pan2023automatically, ; gou2024critic, ). This capability also facilitates a more robust framework, i.e., scalable oversight (bowman2022measuring, ; saunders2022selfcritiquing, ), for ensuring the AI systems growing in scale and capability remain aligned with human-desired outcomes and ethical standards.
So far, while numerous works have been proposed to evaluate critique ability of LLMs in downstream tasks, like common NLP tasks freitag2022results ; fabbri2021summeval and reasoning tasks gou2024critic ; luo2023critique , their comprehensiveness and reliability are limited. Specifically, existing works  (wang2023shepherd, ; li2024generative, ; cui2023ultrafeedback, ) typically evaluate only specific aspects of critique ability, resulting in limited evaluated critique dimensions sun2024critique ; li2024generative ; cui2023ultrafeedback ; wang2023shepherd ; kim2024prometheus , insufficient analysis of response qualities and task types gou2024critic ; luo2023critique . Besides, while GPT-4 is frequently used to evaluate textual or natural language critiques (pan2023automatically, ; wang2023shepherd, ; sun2024critique, ), its reliability across all critique dimensions and tasks remains unverified (wang2023shepherd, ; zhang2024selfcontrast, ). In summary, a comprehensive and reliable benchmark for assessing critique capability of LLMs is still under-explored, significantly impeding the in-depth analysis.
Table 1: Comparison between the test and dev set of benchmarks and CriticEval. A complete list are described in Appendix C. The response quality in some benchmarks is unclassified (-). PR denotes the Pass Rate on reasoning and coding tasks. Our concurrent works are marked with a $\dagger$.
Benchmarks
Critique
Format
Critique
Dimension
Response
Quality
Test NL
Data Size
Subjective
Metric
Objective
Metric
Human
Anno.
Release
CriticBench (luo2023critique, ) Scalar 1 2 0 - Accuracy
Shepherd (wang2023shepherd, ) NL 1 - 352 GPT-4 -
Auto-J (li2024generative, ) NL/Scalar 2 - 232 GPT-4 Accuracy
UltraFeedback (cui2023ultrafeedback, ) NL 1 - 450 GPT-4 -
CriticBench (lin2024criticbench, ) $\dagger$ Scalar 2 2 0 - F1,PR
MetaCritique (sun2024critique, ) $\dagger$ NL 1 - 300 GPT-4 w/ Ref. -
SummEval (fabbri2021summeval, ) Scalar 1 - 0 - Correlation
WMT-22 (zh-en) (freitag2022results, ) Scalar 1 - 0 - Correlation
MT-Bench (zheng2023judging, ) Scalar 1 - 0 - Accuracy
CriticEval (Ours) NL/Scalar 4 4 3,608 GPT-4 w/ Ref.
Correlation,PR
To fill this gap, we propose a novel benchmark, CriticEval, designed to comprehensively and reliably measure critique capability of LLMs. Specifically, to ensure comprehensiveness, CriticEval evaluates critique ability of LLMs from following dimensions: evaluating a single response (feedback), comparing pairs of responses (comparison), correcting the response based on feedback (correction) and evaluating one feedback of LLM (meta-feedback). These critique dimensions cover all categories of critiques in previous works li2024generative ; sun2024critique ; gou2024critic and the necessary capabilities for self-improvement of LLM (yuan2024selfrewarding, ) and scalable oversight (saunders2022selfcritiquing, ). These critique dimensions are measured under nine diverse task scenarios, including three common NLP tasks, two alignment tasks, and four math reasoning and coding tasks. Moreover, evaluated responses in each critique dimension and each task are collected using various open-source and closed-source LLMs with different capabilities, with human annotation ensuring varied quality levels. Furthermore, since both the scalar-valued and textual formats of critique are commonly used in these scenarios (pan2023automatically, ), CriticEval evaluate the critiques in both formats, equipped with objective (fu2023gptscore, ; alpaca_eval, ) and subjective (wang2023shepherd, ; li2024generative, ) evaluations, respectively. Note that scalar-valued critiques typically refer to Likert scores and preference labels, while textual critiques refer to more fine-grained textual analysis about response quality (li2024generative, ; pan2023automatically, ). Overall, as shown in Table 1, CriticEval exhibits significant advantages in comprehensiveness compared to previous benchmarks. It demonstrates great diversity in critique formats, critique dimensions, response qualities and the data size of textual critique.
@@
To ensure the reliability of evaluating textual critiques in CriticEval, a large number of high-quality critiques are annotated, serving as references for GPT-4 to evaluate textual critiques automatically. To annotate these textual critiques efficiently, we employ a human-in-the-loop pipeline (liu2023alignbench, ), first generated by GPT-4 and then rigorously reviewed and refined by human experts.
Extensive evaluations of 35 widely used open-source and closed-source LLMs prove the reliability of CriticEval. Specifically, GPT-4 with human-annotated reference critiques achieves close correlations with human judgments, while removing them results in significant performance loss. Additionally, critiques with higher scores consistently lead to superior improvements, illustrating a clear correlation between the real critique ability of LLMs and their evaluation scores within CriticEval. Then, extensive evaluations results also demonstrate that some open-source LLMs, such as Qwen (qwen, ) and InternLM2 (cai2024internlm2, ), are approaching state-of-the-art closed-source LLMs in critique capabilities, and their critique ability could be further improved through scaling strategy. Besides, the effectiveness of critique datasets is also validated. Finally, these evaluation results also reveal several intriguing phenomena:
  • Critique difficulty varies by task type. For instance, math reasoning and coding tasks are more challenging for feedback, comparison and the self-improvement of LLMs, while they are easier for meta-feedback.
  • There is an inverse relationship between the quality of critiques and responses. For example, high-quality responses pose a greater challenge to critique effectively.
@@
  • Critique difficulty correlates with the critique dimensions; notably, comparison and meta-feedback dimensions present greater challenges than feedback dimension.
@@
These observations and phenomena promote an in-depth understanding of critique ability of LLMs. We hope the discoveries could spur future research in this field.
2 Related Work
2.1 Application of Critique Ability
Automatic Evaluation
Automatic evaluation, also known as critique ability in recent works (pan2023automatically, ; wang2023shepherd, ), has been well studied in the past few years 10.1145/3423168 ; fu2023gptscore . It aims to accurately judge the evaluated responses in numerous NLP tasks and reduce the high cost of human annotations (fu2023gptscore, ; zheng2023judging, ; xu-etal-2023-instructscore, ; li2024dissecting, ; wang2023pandalm, ). Recently, advanced LLMs, like GPT-4, have exhibited very close correlation with human judgments (fu2023gptscore, ; liu2023geval, ; sun2024critique, ), assign textual critiques with corresponding quality scores, i.e., the scalar-valued score, in a chain-of-thought inference manner (wei2023chainofthought, ). To further mitigate the high inference cost of closed-source LLMs, numerous works propose to improve critique ability of open-source LLMs by fine-tuning them on critique datasets generated by GPT-4 (kim2024prometheus, ; ke2023critiquellm, ; jiang2023tigerscore, ), like Auto-J (li2024generative, ) and UltraFeedback (cui2023ultrafeedback, ).
LLM Self-improvement
So far, critique ability has been widely used for self-improvement of LLMs in two stages: (1) Inference stage: Given textual critiques that analyze the flaws in the response and provide suggestions, LLMs can iteratively improve the response quality (saunders2022selfcritiquing, ; zhang2024selfcontrast, ; madaan2023selfrefine, ; yao2023react, ; fernandes-etal-2023-devil, ).; (2) Training stage: Scalar-valued critiques are frequently used to compile responses with a clear performance gap for rejective fine-tuning (RFT) or preference learning (RLHF lee2023rlaif ), which further enhances LLM capabilities (yuan2024selfrewarding, ; xu2023reasons, ; bowman2022measuring, ; bai2022constitutional, ; xu2024chatglmmath, ). For instance, Self-rewarding yuan2024selfrewarding improves Llama-2-70B by fine-tuning it on samples selected based on its rewards. Similarly, ChatGLM-Math xu2024chatglmmath fine-tunes a math-critique model for scoring generated answers, which are used for rejective fine-tuning touvron2023llama and direct preference optimization rafailov2023direct .
2.2 Benchmarking Critique Ability
So far, numerous meta-evaluation benchmarks have been proposed to evaluate the critique ability of models fu2023gptscore . Early benchmarks mainly focus on evaluating scalar-valued critiques (jiang2023tigerscore, ; xu-etal-2023-instructscore, ) on common NLP tasks fu2023gptscore ; 10.1145/3423168 , like translation freitag2022results and summary fabbri2021summeval by computing the correlations between model and human judgments. Recent works also assess scalar-valued critiques of LLMs on reasoning and coding tasks xu-etal-2023-instructscore ; luo2023critique . For example, CriticBench (luo2023critique, ) built from 3 reasoning tasks, analyzes important properties of critique ability of LLMs. Our concurrent work, CriticBench (gou2024critic, ) analyzes several intriguing findings among generation, critique and correction capability on responses collect from five reasoning tasks. [^1] This work has the same name as the previous CriticBench (luo2023critique, ). Compared with these existing works, our proposed CriticEval exhibits advantages on several crucial factors, like critique dimensions, response qualities and diverse task types, leading to more comprehensive evaluations for critique ability. Although CriticBench (luo2023critique, ) collect high-quality responses, their quality levels are still unclassified.
Beyond scalar-valued critiques, evaluating textual critiques is more challenging (wang2023shepherd, ; zhang2024selfcontrast, ). Most existing works coarsely evaluated textual critiques using GPT-4 li2024generative ; cui2023ultrafeedback , proven unreliable (wang2023shepherd, ; sun2024critique, ). Unlike them, our extensive results prove that GPT-4 with human-annotated critiques is reliable for evaluating textual critiques. Although our concurrent work, MetaCritique sun2024critique , demonstrates the reliability of evaluating textual critiques by verifying their Atomic Information Units, it is unclear whether their conclusions could be extended to more critique dimensions and tasks.
Figure 1: Cases of four critique dimensions. Scalar-valued critiques are scores and preference labels.
@@ 3 Preliminaries
We first formally define the key concepts and their corresponding notions in CriticEval. Figure 1 shows a specific case to understand these concepts.
@@
represent the queries and generations of LLMs, respectively.
@@
aims to analyze and refine the generated responses. Formally, this paper studies the critique capabilities in four dimensions: (1) feedback $F_{s}$ involves textual analysis and a quality score. Good feedback should not only find flaws in responses but also provide helpful suggestions for correction (saunders2022selfcritiquing, ); (2) correction or refinement $CR$ aims to revise responses with or without feedback. Previous evaluations (li2024generative, ; cui2023ultrafeedback, ; luo2023critique, ) overlook this dimension, although it is an inevitable step when letting the model improve itself (bai2022constitutional, ); (3) comparison $F_{c}$ contains a textual critique and a preference label for a pair of responses $(R_{a},R_{b})$; (4) Meta-feedback $F_{s}(F_{s})$, i.e., the feedback of feedback itself (saunders2022selfcritiquing, ), involves a rating score reflecting the quality of $F_{s}$ and corresponding textual analysis, which is a high-level critique dimension. Such an ability is necessary to improve critique ability of LLMs (yuan2024selfrewarding, ; sun2024critique, ). Due to the complexity of the meta-feedback dimension, textual critiques are not collected in this paper, and we leave it for future research.
To the best of our knowledge, these four critiques cover all categories of critiques examined in prior research li2024generative ; gou2024critic ; sun2024critique . Although the feedback of correction and comparison dimensions are also important, they are not essential for the self-improvement of LLMs. Thus, this study mainly focuses on studying the feedback of feedback $F_{s}(F_{s})$ and leaving the rest of them for our future work.
4 CriticEval Construction
@@
Given the challenge of crafting scalar-valued and textual critiques from scratch, we construct CriticEval using a human-in-the-loop data construction pipeline as shown in Figure 2.
4.1 Task Input Collection
@@
Task inputs for $9$ distinct tasks are collected to evaluate critique capabilities comprehensively (Step 1 in Figure 2). Specifically, CriticEval includes three widely used tasks for evaluating critique ability: (1) representative classical language tasks: summary (stienon2020learning, ), translation (specia-etal-2020-findings-wmt, ), and question-answering (OpenBookQA2018, ); (2) LLM alignment: general chat scenarios (alpaca_eval, ) and harmlessness cases (bai2022constitutional, ); (3) reasoning and code capabilities: math reasoning with chain-of-thought (CoT) and program-of-thought (PoT), and coding with and without execution results. We hereinafter refer to “code w/ execution” as “CodeExec” and “code w/o execution” as “CodeNE”. For each task, we collect around 100 task inputs from the test sets of some widely used benchmark datasets to ensure the task input quality and avoid data contamination. Please refer to Appendix D for more details about the data source.
4.2 Response and Critique to be Evaluated
For each collected $I$ in each task, LLMs of different scales and capabilities are first employed to generate responses with diverse flaws (Step 2 (a) in Figure 2). The complete list of LLMs is in Appendix E. Then, low-, medium-, and high-quality responses with diverse quality differences are collected according to the quality score annotated by the human raters with GPT-4-turbo as the assistant (Step 2 (b)). Moreover, we also collect golden or correct responses, which have been proven challenging for critiques (zhang2024selfcontrast, ). More details about how to select low-, medium, high-quality and correct responses can be found in Appendix F.
@@ 4.2 Response and Critique to be Evaluated
After collecting responses, we further collect critiques to be evaluated for the meta-feedback dimension by utilizing four LLMs that are known powerful for critiques (Step 3 (d) in Figure 2): (1) GPT-4; (2) GPT-3.5-turbo; (3) Auto-J-13B (li2024generative, ); (4) UltraCM-13B (cui2023ultrafeedback, ).
Figure 2: The data construction pipeline for our proposed CriticEval. Step 1: $9$ tasks and numerous LLMs are prepared. Step 2: LLMs are employed to generate responses, which are then meticulously reviewed by human experts. Step 3: Critiques are generated by LLMs with strong critique ability, and human experts annotate them.
4.3 Reference Critique Generation and Annotation